A positive point charge (q = +5.00 times 10-8 C) i
Answer #1

given,

q = 5 * 10^-8 C

ra = 1.72 m

Wab = -7.47 * 10^-9 J

potential at ra = kq/ra

potential at rb = kq/rb

work done = -E * distance

work done = q0 * (V_final -V_initial)

-7.47 * 10^-9 = -9 * 10^9 * 5 * 10^-8 * ( 1 / rb - 1 / 1.72) * 3.33 * 10^-11

rb = 0.926 m